OpenStudy (anonymous):

solve for K: log base 4 (5k+4)=3

OpenStudy (anonymous):

20k+16=3

OpenStudy (anonymous):

20+16=36

OpenStudy (anonymous):

can you show me the steps? i

OpenStudy (anonymous):

3/36=12

OpenStudy (anonymous):

k=12

OpenStudy (anonymous):

i need the steps/// not the answer. thanks though. i am trying to learn how to do this

OpenStudy (anonymous):

first u do the distributive property

OpenStudy (anonymous):

LOG base 4..... does what to the problem??

OpenStudy (anonymous):

u distribute that 4 to all of the numbers

OpenStudy (anonymous):

is that what the base 4 does?

OpenStudy (anonymous):

yeah u distrbute it